Daniel plans to visit Australia. Whether he goes alone or with a companion is equally likely. If he travels with a companion there is a 40% chance of joining a guided tour. If he travels alone, there is an 80% chance of joining a guided tour. A) What is the probability of Daniel travelling with a companion and not joining a guided tour? B) What is the probability of Daniel joining a guided tour?
Answer by ikleyn(52906)

.

(A)  0.5*(1 - 0.4) = 0.5*0.6 = 0.3 = 30%.          ANSWER


(B)  0.5*0.4 + 0.5*0.8 = 0.2 + 0.4 = 0.6 = 60%.    ANSWER

Answered and solved.